Melissa takes classes at Westside community college and Pinewood community college. At Westside, class fees are $98 per credit hour, and at Pinewood class fees are $115 per credit hour. Melissa is taking a combined total of 19 credit hours at the two schools .
Suppose that she is taking W credit hours at Westside. Right and expression for the combined total dollar amount she paid for her class fees.

Answer by Theo(13342)

if she is taking w credit hours at westside, then she is paying 98 * w for the credit hours at westside.

is she is taking p credit hours at pinewood, then she is paying 115 * p for the credit hours at pinewood.

the total she is paying for classes at westside and pinewood would be 98 * w + 115 * p

the total credits she is taking is 19, therefore w + p = 19.

since you know that w + p = 19, then you could solve for p to get p = 19 - w.

you could then replace p in the total cost equation with 19 - w to get total cost = 98 * w + 115 * (19 - w)